OpenStudy (anonymous):

31-8(y+5)=7 help me solf this

OpenStudy (anonymous):

Easy, Your equation look's like this. \[31 - 8 ( y + 5 ) = 7 \] Step 1 : Take 8 adding the y putting it together looking like this \[8y - 16 = 0\] Step 2 : \[-8y - 9 = 7\] Simplify and your answer should look like this y = -2 Hope this helps!

OpenStudy (anonymous):

thank youu! yes it did